a) Frenzi Communications Inc. (FCI) has the following capital structure, which it considers to be optimal: debt = 25%, preferred stock = 15%, and common stock = 60%. FCI’s tax rate is 40%, and investors expect earnings and dividends to grow at a constant rate of 6% in the future. FCI paid a dividend of Gh₵3.70 per share last year (D0), and its stock currently sells at a price of Gh₵60 per share. Ten-year Treasury bonds yield 6%, the market risk premium is 5%, and FCI’s beta is 1.3. The following terms would apply to new security offerings. Preferred: New preferred could be sold to the public at a price of Gh₵100 per share, with a dividend of Gh₵9. Flotation costs of Gh₵5 per share would be incurred. Debt: Debt could be sold at an interest rate of 9%. Common: New common equity will be raised only by retaining earnings. i. Find the component costs of debt, preferred stock, and common stock. (9 marks) ii. What is the WACC? b) Identify and describe the phases of capital budgeting.


1
Expert's answer
2021-09-30T12:20:53-0400

i. The component costs are:

Cost of debt is:

Kd = 0.09.

Cost of preferred stock is:

"Kp = \frac{9}{100-5} = 0.0947."

Cost of common stock is:

Ke = 0.06 + 1.3*(0.05 - 0.06) = 0.047.

ii. "WACC = 0.25*0.09*(1 - 0.4) + 0.15*0.0947 + 0.6*0.047 = 0.0559."

b) The capital budgeting process consists of five steps:

  • Identify and evaluate potential opportunities. 
  • Estimate operating and implementation costs.
  • Estimate cash flow or benefit.
  • Assess risk.
  • Implement.